OpenStudy (anonymous):

Simplify the following √27-√3

OpenStudy (anonymous):

\[\sqrt{3^3} -\sqrt3\] \[\sqrt3(3 - 1 )\] \[2 \sqrt3\]

OpenStudy (mimi_x3):

3 sqaure root 3 - sqaure root 3 = 2 square root 3

OpenStudy (mimi_x3):

argh how fast xD

OpenStudy (gg):

\[\sqrt{27}-\sqrt{3}=\sqrt{9*3}-\sqrt{3}=\sqrt{9}\sqrt{3}-\sqrt{3}=3\sqrt{3}-\sqrt{3}=\sqrt{3}(3-1)=2\sqrt{3}\]

OpenStudy (anonymous):

ok that's good but type the value too i.e, 3.464

OpenStudy (anonymous):

thankyoouu :)

OpenStudy (anonymous):

\[\sqrt{3^2*3} -\sqrt{3}\] \[3\sqrt{3}-\sqrt{3}\] take \[\sqrt{3}\] as common \[\sqrt{3}(3-1)\] \[2\sqrt{3}\]

OpenStudy (anonymous):

I believe I did that earlier

OpenStudy (anonymous):

do u have any problem Ishaan that i had answered this question

OpenStudy (mimi_x3):

LOL, he has xD

OpenStudy (anonymous):

yea now it seems to be like that :)

OpenStudy (mimi_x3):

lols i was jk

OpenStudy (anonymous):

Of course I have problem with that kush

OpenStudy (anonymous):

c'mon fellas chill

OpenStudy (anonymous):

that's nice ishaan , but why ?

OpenStudy (mimi_x3):

Because his mean :P jkss

OpenStudy (anonymous):

ah i don't wanna chill it's bee a while since I have had fight with someone ; ) lol

OpenStudy (anonymous):

ah MIMI is always right I can't prove her wrong I am mean

OpenStudy (anonymous):

ah me too ishaan , let us fight , who is going to start that u or me :) haha

OpenStudy (mimi_x3):

LOL, yup im always right not like Pink Knight who's always wrong xDD

OpenStudy (anonymous):

lol not Pink Knight \[\Huge{\textbf{Dark Knight}}\]

OpenStudy (anonymous):

how u did it ishaan ?

OpenStudy (anonymous):

silent guardian, a watchful protector ; )

OpenStudy (anonymous):

omg pl. stop all this

OpenStudy (mimi_x3):

Well, you PINK , so your a PINK KNIGHT too bad your not DARK

OpenStudy (anonymous):

i'm getiin too much disturbed

OpenStudy (mimi_x3):

LOL, sorry not my fault. Its Pink Knight's xD

OpenStudy (anonymous):

what makes you think I am gonna tell you kush lol well seriously you need to learn LateX only then I can tell you

OpenStudy (anonymous):

ah okay I won't do it again this is my last reply

OpenStudy (mimi_x3):

LOL, see~ how mean ishaan is :Pjks

OpenStudy (anonymous):

ah I lied this is my last reply Mimi is right I am very mean ...bad ishaan bad

OpenStudy (mimi_x3):

Ishaan's a LIAR xD jks

OpenStudy (anonymous):

okay I lied again this should be my last reply ....ah mimi is right again I am a LIar

OpenStudy (mimi_x3):

lolss , you alwayss lie xD We should stop now xD Incase the person gets pissed off xD

OpenStudy (anonymous):

*stopped*

OpenStudy (mimi_x3):

LOL, impossible for ishaan aka PINK KNIGHT :P